OpenStudy (anonymous):

if a uniform chain of length "l"slips from a chair nd it1/3part slips frm chair ..can i knw hw much kinetic energy it vill take to slip frm the chair...

OpenStudy (mos1635):

poore frasing please elaborate....

OpenStudy (anonymous):

u hv just viewd my quest....u knw the answer...

OpenStudy (mos1635):

no iam sorry i do not i am not even sure i have understood correctly what exactly the question is. my appologies.

OpenStudy (anonymous):

itz out of reach.....u cant understand it...

OpenStudy (mos1635):

 to be sliping must weirth of hanging chain > friction of rest

OpenStudy (mos1635):

if so Velocity mast be calculated dy deferencial equation due to variable to M hanging and friction

OpenStudy (anonymous):

IN SOLVING KINETIC ENERGY QUESTION WE MUST HAVE MASS AND THE VELOCITY OF THE BODY. WHAT ARE THEY FROM YOUR QUESTION ?

OpenStudy (mos1635):

feel free to help me understand it

OpenStudy (anonymous):

ohh srry mass dena bool gaya...mass is 'm '...nw calculate

OpenStudy (anonymous):

nd velocity "v"

OpenStudy (mos1635):

BUT if you have been given that the chain is sliping with a steady speed you could calculate the kinetic energy of hanging part

OpenStudy (anonymous):

nw i have given all data.....mass,velocita nd length....

OpenStudy (mos1635):

is that velocity steady or it only coresponds to time that 1/3 of chain is hanging?

OpenStudy (anonymous):

only correspond to tym....

OpenStudy (mos1635):

 preservation of energy (assuming no friction) 2=all chain hang 1=1/3 of chain hang K1+U1=K2+U2 1/2 M V1^2 +(-(M/3)*g*(L/6)=1/2 M V2^2+ (-M*g*(L/2) this should give you final velocity.... i am not sure what :hw much kinetic energy it vill take